Grade 12Physical Chemistry

How many mL of each of two hydrochloric acids of strength 12N and 13N are to be mixed to make one litre of 6N solution?

ragini kumari
8 Years agoGrade 12

1 Answer

Abhiraj Pathak
8 Years ago
Let x be the volume of 12N HCl.
and y be the volume of 13N HCl
V1 + V2 = 1L = 1000mL ; 
x + y = 1000L
Since this is a dilution of acids, the number of moles will remain same.
We can use; (N1V1+N2V2)/(V1+V2) = N3
N1 =12N; N2 = 13N; N3 = 6N
12x + 13y = 6000
Substituting x = 1000-y
we get y = 240mL and x = 760mL